OpenStudy (anonymous):

Enter the equation of the circle described below. Center (5, 2), radius = 3

OpenStudy (anonymous):

if you know how to do these and don't mind helping with more. Enter the equation of the circle described below. Center (0, 4), radius Enter the equation of the circle described below. Center (-3, 0), radius Enter the equation of the circle described below. Center (0, 0), radius = 4

OpenStudy (anonymous):

they are all the same if the center is \((h,k)\) and the radius is \(r\) then the equation is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

for example Center (5, 2), radius = 3 equation is \[(x-5)^2+(y-2)^2=9\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

yw